# A unified relative entropy framework for macroscopic limits of Vlasov–Fokker–Planck equations

### v1

@@ 4.1. Diffusive limit
$\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\varphi(x,\xi)(f^{\varepsilon}-M_{\rho,0})\,dxd\xi$
$\quad=\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\varphi(x,\xi)(f^{\varepsilon}-M_{\rho,{\rm u_{\varepsilon}}})\,dxd\xi+\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\varphi(x,\xi)(M_{\rho,{\rm u_{\varepsilon}}}-M_{\rho,0})\,dxd\xi$
$\quad=:I+II,$
where it is clear that
$|I|\leq C{\rm d}_{\rm BL}(f^{\varepsilon},M_{\rho,{\rm u_{\varepsilon}}}),$
and by using a similar argument as in the proof of Lemma 3.2
$|II|\leq C\|\rho\|_{L^{1}}\|{\rm u_{\varepsilon}}\|_{L^{\infty}}\leq C\varepsilon.$
@@
${\rm d}^{2}_{\rm BL}(f^{\varepsilon},M_{\rho,0})\leq C{\rm d}^{2}_{\rm BL}(f^{\varepsilon},M_{\rho,{\rm u_{\varepsilon}}})+C\varepsilon^{2},$
and subsequently,
$\int_{0}^{T}{\rm d}^{2}_{\rm BL}(f^{\varepsilon},M_{\rho,0})\,dt\leq C{\rm d}^{2}_{\rm BL}(\rho^{\varepsilon}_{0},\rho_{0})+C\varepsilon^{2}\left(\mathscr{H}[f^{\varepsilon}|M_{\rho,0}](0)+\mathscr{P}[\rho^{\varepsilon}|\rho](0)\right)+C\varepsilon^{2}.$
This provides the convergence of $\rho^{\varepsilon}$ and $f^{\varepsilon}$ in the BL topology, which implies in particular the weak-* convergence
$f^{\varepsilon}\overset{*}{\rightharpoonup}M_{\rho,0}\quad\mbox{in }L^{2}(0,T;\mathcal{M}(\mathbb{R}^{d}\times\mathbb{R}^{d})),\quad\int_{\mathbb{R}^{d}}f^{\varepsilon}\,d\xi\overset{*}{\rightharpoonup}\rho\quad\mbox{in }L^{\infty}(0,T;\mathcal{M}(\mathbb{R}^{d})).$
Finally, we turn to the momentum variable $\frac{1}{\varepsilon}\rho^{\varepsilon}u^{\varepsilon}$. The dissipation bound in the relative entropy estimate (4.6) gives
$\frac{1}{\varepsilon^{2}}\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho^{\varepsilon}|u^{\varepsilon}|^{2}\,dxdt\leq\frac{2}{\varepsilon^{2}}\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho^{\varepsilon}|u^{\varepsilon}-{\rm u_{\varepsilon}}|^{2}\,dxdt+\frac{2T}{\varepsilon^{2}}\|{\rm u_{\varepsilon}}\|_{L^{\infty}}^{2}\leq C,$
@@
$\int_{0}^{T}\int_{\mathbb{R}^{d}}\frac{1}{\varepsilon}|\rho^{\varepsilon}u^{\varepsilon}|\,dxdt\leq\left(\frac{1}{\varepsilon^{2}}\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho^{\varepsilon}|u^{\varepsilon}|^{2}\,dxdt\right)^{\frac{1}{2}}\leq C$
for some $C>0$ independent of $\varepsilon$. Thus, by Banach–Alaoglu theorem, there exists $\tilde{u}$ such that
$\frac{1}{\varepsilon}\rho^{\varepsilon}u^{\varepsilon}\overset{*}{\rightharpoonup}\tilde{u}\quad\mbox{in }\mathcal{M}((0,T)\times\mathbb{R}^{d}).$ (4.7)
To identify the weak limit $\tilde{u}$, we observe from (4.1) that for any $\varphi\in C_{c}^{\infty}([0,T)\times\mathbb{R}^{d})$
$\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(\rho^{\varepsilon}\partial_{t}\varphi+\frac{1}{\varepsilon}\rho^{\varepsilon}u^{\varepsilon}\cdot\nabla\varphi\right)dxdt+\int_{\mathbb{R}^{d}}\rho^{\varepsilon}_{0}\varphi(x,0)\,dx.$
We then use $\rho^{\varepsilon}\overset{*}{\rightharpoonup}\rho$ in $L^{\infty}(0,T;\mathcal{M}(\mathbb{R}^{d}))$ and (4.7) to obtain
$\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(\rho\partial_{t}\varphi+\tilde{u}\cdot\nabla\varphi\right)dxdt+\int_{\mathbb{R}^{d}}\rho_{0}\varphi(x,0)\,dx.$
On the other hand, $\rho$ satisfies
$\int_{0}^{T}\int_{\mathbb{R}^{d}}\left(\rho\partial_{t}\varphi-\rho(\nabla(-\Delta)^{-\alpha}\rho+\nabla\log\rho)\cdot\nabla\varphi\right)dxdt+\int_{\mathbb{R}^{d}}\rho_{0}\varphi(x,0)\,dx,$
and this implies
[…]
Hence we have
[…]
and subsequently
$\frac{1}{\varepsilon}\rho^{\varepsilon}u^{\varepsilon}\overset{*}{\rightharpoonup}-\rho(\nabla(-\Delta)^{-\alpha}\rho+\nabla\log\rho)\quad\mbox{in }\mathcal{M}((0,T)\times\mathbb{R}^{d}).$
This completes the proof.

### v2

@@ 4.1. Diffusive limit
$\qquad-\varepsilon\int_{0}^{t}\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\Delta\mathcal{H}_{t-s}(x-y)(\rho^{\varepsilon}u^{\varepsilon})(s,y)\cdot\nabla\varphi(x)\,dydxds-\varepsilon\int_{\mathbb{R}^{d}}\rho^{\varepsilon}u^{\varepsilon}(t,x)\cdot\nabla\varphi\,dx$
$\qquad+\varepsilon\int_{\mathbb{R}^{d}}\mathcal{H}_{t}\star_{x}(\rho_{0}^{\varepsilon}u_{0}^{\varepsilon})\cdot\nabla\varphi\,dx+\int_{0}^{t}\int_{\mathbb{R}^{d}}\left(\int_{\mathbb{R}^{d}}(\xi\otimes\xi-\mathbb{I}_{d})f^{\varepsilon}\,d\xi\right):\mathcal{H}_{t-s}\star_{x}\nabla\otimes\nabla\varphi\,dxds$
$\quad=:\sum_{i=1}^{7}I_{i}.$
We estimate the terms $I_{i}$ one by one. To begin with, one readily obtains
$I_{1}\leq\|\rho_{0}^{\varepsilon}-\rho_{0}\|_{H^{-\beta}}\|\mathcal{H}_{t}\star\varphi\|_{H^{\beta}}\leq C\|\varphi\|_{H^{\beta}}\|\rho_{0}^{\varepsilon}-\rho_{0}\|_{H^{-\beta}},$
and
$I_{2}$ $\leq\int_{0}^{t}\|(\rho^{\varepsilon}-\rho)(s)\|_{\dot{H}^{-\alpha}}^{2}\|\mathcal{H}_{t-s}\star_{x}\nabla\varphi\|_{W_{x}^{1,\infty}}\,ds$
$\leq C\left(\mathscr{H}[f^{\varepsilon}|M_{\rho,0}](0)+\mathscr{P}[\rho^{\varepsilon}|\rho](0)+\varepsilon^{2}\right)\int_{0}^{t}\|\mathcal{H}_{t-s}\star_{x}\nabla\varphi\|_{H^{\beta}}\,ds$
$\leq C\|\varphi\|_{H^{\beta}}\left(\mathscr{H}[f^{\varepsilon}|M_{\rho,0}](0)+\mathscr{P}[\rho^{\varepsilon}|\rho](0)+\varepsilon^{2}\right)\int_{0}^{t}(t-s)^{-\frac{1}{2}}\,ds$
$\leq C\|\varphi\|_{H^{\beta}}\left(\mathscr{H}[f^{\varepsilon}|M_{\rho,0}](0)+\mathscr{P}[\rho^{\varepsilon}|\rho](0)+\varepsilon^{2}\right),$
where we used
$x^{k}e^{-ax}\leq\left(\frac{k}{a}\right)^{k}e^{-k}\quad\mbox{on}\quad x>0$
and
$\|\mathcal{H}_{t-s}\star\nabla\varphi\|_{H^{\beta}}^{2}\leq\int_{\mathbb{R}^{d}}(1+|\eta|^{2})^{\beta}|\eta|^{2}e^{-2(t-s)|\eta|^{2}}|\hat{\varphi}|^{2}\,d\eta\leq C(t-s)^{-1}\|\varphi\|_{H^{\beta}}^{2}.$
Next, we decompose $I_{3}$ as
$I_{3}$ $=-\int_{0}^{t}\int_{\mathbb{R}^{d}}(\rho^{\varepsilon}-\rho)(s)(\nabla(-\Delta)^{-\alpha}\rho)(s)\cdot(\mathcal{H}_{t-s}\star_{x}\nabla\varphi)\,dxds$
$\quad+\int_{0}^{t}\int_{\mathbb{R}^{d}}(\rho^{\varepsilon}-\rho)(s)\nabla\cdot(-\Delta)^{-\alpha}(\rho(s)\mathcal{H}_{t-s}\star_{x}\nabla\varphi)\,dxds$
$=:I_{31}+I_{32}.$
For the first part $I_{31}$, we use
$\|f\|_{H^{\beta}}\leq C(\|f\|_{L^{2}}+\|f\|_{\dot{H}^{\beta}}),\quad\|uv\|_{\dot{H}^{\beta}}\leq C(\|u\|_{L^{\infty}}\|v\|_{\dot{H}^{\beta}}+\|u\|_{\dot{H}^{\beta}}\|v\|_{L^{\infty}})$ (4.8)
to have
$I_{31}$ $\leq\int_{0}^{t}\|\rho^{\varepsilon}-\rho\|_{H^{-\beta}}\|\nabla(-\Delta)^{-\alpha}\rho(s)\cdot(\mathcal{H}_{t-s}\star_{x}\nabla\varphi)\|_{H^{\beta}}\,ds$
$\leq C\int_{0}^{t}\|\rho^{\varepsilon}-\rho\|_{H^{-\beta}}\|\nabla(-\Delta)^{-\alpha}\rho(s)\cdot(\mathcal{H}_{t-s}\star_{x}\nabla\varphi)\|_{L^{2}}\,ds$
$\quad+C\int_{0}^{t}\|\rho^{\varepsilon}-\rho\|_{H^{-\beta}}\|\nabla(-\Delta)^{-\alpha}\rho(s)\cdot(\mathcal{H}_{t-s}\star_{x}\nabla\varphi)\|_{\dot{H}^{\beta}}\,ds$
$\leq C\|\nabla(-\Delta)^{-\alpha}\rho\|_{L^{\infty}(0,T;L^{\infty}\cap\dot{H}^{\beta})}\int_{0}^{t}\|\rho^{\varepsilon}-\rho\|_{H^{-\beta}}\|\mathcal{H}_{t-s}\star_{x}\nabla\varphi\|_{H^{\beta}}\,ds$
$\leq C\|\nabla(-\Delta)^{-\alpha}\rho\|_{L^{\infty}(0,T;L^{\infty}\cap\dot{H}^{\beta})}\|\varphi\|_{H^{\beta}}\int_{0}^{t}(t-s)^{-\frac{1}{2}}\|\rho^{\varepsilon}-\rho\|_{H^{-\beta}}\,ds.$
For the second part $I_{32}$, we distinguish the cases $\alpha\in(\frac{1}{2},1)$ and $\alpha\in(0,\frac{1}{2}]$. If $\alpha\in(\frac{1}{2},1)$, the Hardy–Littlewood–Sobolev inequality yields
$\|\nabla\cdot(-\Delta)^{-\alpha}(\rho(s)\mathcal{H}_{t-s}\star_{x}\nabla\varphi)\|_{L^{2}}$ $\leq C\|\rho(s)\mathcal{H}_{t-s}\star_{x}\nabla\varphi\|_{\dot{H}^{1-2\alpha}}$
$\leq C\|\rho(s)\mathcal{H}_{t-s}\star_{x}\nabla\varphi\|_{L^{p}}$
$\leq C\|\rho\|_{L^{\infty}(0,T;L^{p})}\|\mathcal{H}_{t-s}\star_{x}\nabla\varphi\|_{L^{\infty}}$
$\leq C\|\rho\|_{L^{\infty}(0,T;L^{p})}\|\mathcal{H}_{t-s}\star_{x}\varphi\|_{H^{\beta}},$
where $p>1$ satisfies $\frac{1}{2}=\frac{1}{p}-\frac{2\alpha-1}{d}$. If $\alpha\in(0,\frac{1}{2}]$, one simply gets
$\|\nabla\cdot(-\Delta)^{-\alpha}(\rho(s)\mathcal{H}_{t-s}\star_{x}\nabla\varphi)\|_{L^{2}}$
$\quad\leq C\|\rho(s)\mathcal{H}_{t-s}\star_{x}\nabla\varphi\|_{\dot{H}^{1-2\alpha}}$
$\quad\leq C\left(\|\rho\|_{L^{\infty}(0,T;L^{\infty})}\|\mathcal{H}_{t-s}\star_{x}\nabla\varphi\|_{\dot{H}^{1-2\alpha}}+\|\rho\|_{L^{\infty}(0,T;\dot{H}^{1-2\alpha})}\|\mathcal{H}_{t-s}\star_{x}\nabla\varphi\|_{L^{\infty}}\right)$
$\quad\leq C\|\rho\|_{L^{\infty}(0,T;H^{\beta})}\|\mathcal{H}_{t-s}\star_{x}\varphi\|_{H^{\beta}}.$
Combining these bounds with (4.8), we get
$I_{32}$ $\leq\int_{0}^{t}\|\rho^{\varepsilon}-\rho\|_{H^{-\beta}}\|\nabla\cdot(-\Delta)^{-\alpha}(\rho(s)\mathcal{H}_{t-s}\star_{x}\nabla\varphi)\|_{H^{\beta}}\,ds$
$\leq C\int_{0}^{t}\|\rho^{\varepsilon}-\rho\|_{H^{-\beta}}\|\nabla\cdot(-\Delta)^{-\alpha}(\rho(s)\mathcal{H}_{t-s}\star_{x}\nabla\varphi)\|_{L^{2}}\,ds$
$\quad+C\int_{0}^{t}\|\rho^{\varepsilon}-\rho\|_{H^{-\beta}}\|\rho(s)\mathcal{H}_{t-s}\star_{x}\nabla\varphi\|_{\dot{H}^{\beta+1-2\alpha}}\,ds$
$\leq C\|\rho\|_{L^{\infty}(0,T;L^{p}\cap H^{\beta})}\int_{0}^{t}\|\rho^{\varepsilon}-\rho\|_{H^{-\beta}}\|\mathcal{H}_{t-s}\star_{x}\varphi\|_{H^{\beta}}\,ds$
$\quad+C\|\rho\|_{L^{\infty}(0,T;L^{\infty}\cap H^{\beta+1-2\alpha})}\int_{0}^{t}\|\rho^{\varepsilon}-\rho\|_{H^{-\beta}}\|\mathcal{H}_{t-s}\star_{x}\varphi\|_{H^{\beta+2-2\alpha}}\,ds$
$\leq C\|\rho\|_{L^{\infty}(0,T;L^{p}\cap H^{\beta+1})}\|\varphi\|_{H^{\beta}}\int_{0}^{t}(t-s)^{-1+\alpha}\|\rho^{\varepsilon}-\rho\|_{H^{-\beta}}\,ds.$
We now turn to the remaining terms. For $I_{4}$, we use $\int_{\mathbb{R}^{d}}\rho^{\varepsilon}|u^{\varepsilon}-{\rm u}_{\varepsilon}|^{2}\,dx\leq\mathscr{H}[f^{\varepsilon}|M_{\rho,{\rm u}_{\varepsilon}}]$ to obtain
$I_{4}$ $=-\varepsilon\int_{0}^{t}\int_{\mathbb{R}^{d}}\rho^{\varepsilon}u^{\varepsilon}\mathcal{H}_{t-s}\star_{x}(\Delta\nabla\varphi)\,dxds$
$\leq C\varepsilon\|\varphi\|_{H^{\beta}}\int_{0}^{t}\max\left\{1,(t-s)^{-\frac{\frac{d}{2}+3-\beta}{2}}\right\}\left(\|\rho^{\varepsilon}(u^{\varepsilon}-{\rm u}_{\varepsilon})\|_{L^{1}}+\|\rho^{\varepsilon}{\rm u}_{\varepsilon}\|_{L^{1}}\right)\,ds$
$\leq C\varepsilon\|\varphi\|_{H^{\beta}}\left(\int_{0}^{t}\left[1+(t-s)^{-\frac{\frac{d}{2}+3-\beta}{2}}\right]\,ds\right)^{\frac{1}{2}}\left(\int_{0}^{t}\int_{\mathbb{R}^{d}}(t-s)^{-\frac{\frac{d}{2}+3-\beta}{2}}\rho^{\varepsilon}|u^{\varepsilon}-{\rm u}_{\varepsilon}|^{2}\,dxds\right)^{\frac{1}{2}}+C\varepsilon^{2}\|\varphi\|_{H^{\beta}}$
$\leq C\varepsilon\|\varphi\|_{H^{\beta}}\left(\mathscr{H}[f^{\varepsilon}|M_{\rho,{\rm u}_{\varepsilon}}](0)+\mathscr{P}[\rho^{\varepsilon}|\rho](0)+\varepsilon^{2}\right)^{\frac{1}{2}},$
where we used $\beta>\frac{d}{2}+1$ (i.e. $\frac{\frac{d}{2}+3-\beta}{2}<1$) and
$\|\mathcal{H}_{t-s}\star_{x}\Delta\nabla\varphi\|_{L^{\infty}}$ $\leq C\int_{\mathbb{R}^{d}}|\eta|^{3}e^{-(t-s)|\eta|^{2}}|\hat{\varphi}|\,d\eta$
$\leq C\left(\int_{\mathbb{R}^{d}}\frac{|\eta|^{6}}{(1+|\eta|^{2})^{\beta}}e^{-(t-s)|\eta|^{2}}\,d\eta\right)^{\frac{1}{2}}\|\varphi\|_{H^{\beta}}$
$\leq C\max\left\{1,(t-s)^{-\frac{\frac{d}{2}+3-\beta}{2}}\right\}\|\varphi\|_{H^{\beta}}.$
For $I_{5}$, we find
$I_{5}$ $\leq C\varepsilon\|\rho^{\varepsilon}u^{\varepsilon}\|_{L^{1}}\|\nabla\varphi\|_{L^{\infty}}$
$\leq C\varepsilon\left(\|\rho^{\varepsilon}(u^{\varepsilon}-{\rm u}_{\varepsilon})\|_{L^{1}}+\varepsilon\right)\|\varphi\|_{H^{\beta}}$
$\leq C\varepsilon\|\varphi\|_{H^{\beta}}\left(\mathscr{H}[f^{\varepsilon}|M_{\rho,{\rm u}_{\varepsilon}}](0)+\mathscr{P}[\rho^{\varepsilon}|\rho](0)+\varepsilon^{2}\right)^{\frac{1}{2}}.$
Similarly, for the initial contribution $I_{6}$, we have
$I_{6}$ $\leq\varepsilon\|\rho_{0}^{\varepsilon}u_{0}^{\varepsilon}\|_{L^{1}}\|\mathcal{H}_{t}\star_{x}\nabla\varphi\|_{L^{\infty}}\leq C\varepsilon\|\varphi\|_{H^{\beta}}\left(\mathscr{H}[f^{\varepsilon}|M_{\rho,{\rm u}_{\varepsilon}}](0)+\varepsilon^{2}\right)^{\frac{1}{2}}.$
It remains to estimate $I_{7}$, which contains the second-order velocity moment. To this end, we first observe that
$\int_{\mathbb{R}^{d}}(\xi\otimes\xi-\mathbb{I}_{d})f^{\varepsilon}\,d\xi$ $=\int_{\mathbb{R}^{d}}(\xi\otimes\xi-u^{\varepsilon}\otimes u^{\varepsilon}-\mathbb{I}_{d})f^{\varepsilon}\,d\xi+\rho^{\varepsilon}u^{\varepsilon}\otimes u^{\varepsilon}$
$=\int_{\mathbb{R}^{d}}\left[((\xi-u^{\varepsilon})f^{\varepsilon}+\nabla_{\xi}f^{\varepsilon})\otimes\xi+u^{\varepsilon}\otimes((\xi-u^{\varepsilon})f^{\varepsilon}+\nabla_{\xi}f^{\varepsilon})\right]\,d\xi+\rho^{\varepsilon}u^{\varepsilon}\otimes u^{\varepsilon}.$
@@
$I_{7}$ $\leq\int_{0}^{t}\left(\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\frac{1}{f^{\varepsilon}}|(\xi-u^{\varepsilon})f^{\varepsilon}+\nabla_{\xi}f^{\varepsilon}|^{2}\,dxd\xi\right)^{\frac{1}{2}}\left(\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}|\xi|^{2}f^{\varepsilon}\,dxd\xi\right)^{\frac{1}{2}}\|\mathcal{H}_{t-s}\star_{x}\nabla\otimes\nabla\varphi\|_{L_{x}^{\infty}}\,ds$
$\quad+\int_{0}^{t}\left(\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\frac{1}{f^{\varepsilon}}|(\xi-u^{\varepsilon})f^{\varepsilon}+\nabla_{\xi}f^{\varepsilon}|^{2}\,dxd\xi\right)^{\frac{1}{2}}\left(\int_{\mathbb{R}^{d}}\rho^{\varepsilon}|u^{\varepsilon}|^{2}\,dx\right)^{\frac{1}{2}}\|\mathcal{H}_{t-s}\star_{x}\nabla\otimes\nabla\varphi\|_{L_{x}^{\infty}}\,ds$
$\quad+\int_{0}^{t}\int_{\mathbb{R}^{d}}\left(\rho^{\varepsilon}|u^{\varepsilon}-{\rm u}_{\varepsilon}|^{2}+\rho^{\varepsilon}|{\rm u}_{\varepsilon}|^{2}\right)\|\mathcal{H}_{t-s}\star_{x}\nabla\otimes\nabla\varphi\|_{L_{x}^{\infty}}\,dxds$
$\leq C\|\varphi\|_{H^{\beta}}\left(\int_{0}^{t}(t-s)^{-\frac{1}{2}}\left(\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\frac{1}{f^{\varepsilon}}|(\xi-u^{\varepsilon})f^{\varepsilon}+\nabla_{\xi}f^{\varepsilon}|^{2}(s)\,dxd\xi\right)ds\right)^{\frac{1}{2}}$
$\quad+C\|\varphi\|_{H^{\beta}}\int_{0}^{t}(t-s)^{-\frac{1}{2}}\left(\mathscr{H}[f^{\varepsilon}|M_{\rho,{\rm u}_{\varepsilon}}](0)+\varepsilon^{2}\right)\,dxds$
$\leq C\|\varphi\|_{H^{\beta}}\left(\int_{0}^{t}(t-s)^{-\frac{1}{2}}\left(\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\frac{1}{f^{\varepsilon}}|(\xi-u^{\varepsilon})f^{\varepsilon}+\nabla_{\xi}f^{\varepsilon}|^{2}(s)\,dxd\xi\right)ds\right)^{\frac{1}{2}}$
$\quad+C\|\varphi\|_{H^{\beta}}\left(\mathscr{H}[f^{\varepsilon}|M_{\rho,{\rm u}_{\varepsilon}}](0)+\mathscr{P}[\rho^{\varepsilon}|\rho](0)+\varepsilon^{2}\right).$
Collecting the above estimates, and using (4.6), we arrive at
$\begin{aligned} \|(\rho^{\varepsilon}-\rho)(t)\|_{H^{-\beta}}&\leq C\|\rho_{0}^{\varepsilon}-\rho_{0}\|_{H^{-\beta}}+C\left(\mathscr{H}[f^{\varepsilon}|M_{\rho,0}](0)+\mathscr{P}[\rho^{\varepsilon}|\rho](0)+\varepsilon^{2}\right)\\ &\quad+C\int_{0}^{t}\left((t-s)^{-\frac{1}{2}}+(t-s)^{-1+\alpha}\right)\|(\rho^{\varepsilon}-\rho)(s)\|_{H^{-\beta}}\,ds\\ &\quad+C\left(\int_{0}^{t}(t-s)^{-\frac{1}{2}}\left(\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\frac{1}{f^{\varepsilon}}|(\xi-u^{\varepsilon})f^{\varepsilon}+\nabla_{\xi}f^{\varepsilon}|^{2}(s)\,dxd\xi\right)ds\right)^{\frac{1}{2}}\end{aligned}$ (4.9)
due to (4.6).
To close the above estimate, we use the following auxiliary lemma.
Lemma 4.2.
Let $T>0$, $\alpha>0$, and let $f,g\in L^{2}(0,T)$ be nonnegative functions such that
$f(t)\leq g(t)+C\int_{0}^{t}\bigl((t-s)^{-\frac{1}{2}}+(t-s)^{-1+\alpha}\bigr)f(s)\,ds$
for a.e. $t\in(0,T)$. Then there exists a constant $C_{T}>0$, depending only on $C$, $\alpha$, and $T$, such that
$\|f\|_{L^{2}(0,T)}\leq C_{T}\|g\|_{L^{2}(0,T)}.$
Proof.
Define
$\kappa(t):=C\bigl(t^{-\frac{1}{2}}+t^{-1+\alpha}\bigr),\quad t\in(0,T),$
and, for measurable functions $\phi,\psi$ on $(0,T)$, set
$(\phi\star_{t}\psi)(t):=\int_{0}^{t}\phi(t-s)\psi(s)\,ds.$
Then the assumption reads
$f\leq g+\kappa\star_{t}f\quad\text{a.e. on }(0,T).$
For $\lambda>0$, set
$F(t):=e^{-\lambda t}f(t),\quad G(t):=e^{-\lambda t}g(t),$
and define
$\kappa_{\lambda}(t):=e^{-\lambda t}\kappa(t)=Ce^{-\lambda t}\left(t^{-\frac{1}{2}}+t^{-1+\alpha}\right).$
Then, we obtain
$F(t)\leq G(t)+(\kappa_{\lambda}\star_{t}F)(t).$
Since $\alpha>0$, we find $\kappa_{\lambda}\in L^{1}(0,T)$ and
$\|\kappa_{\lambda}\|_{L^{1}(0,T)}=C\int_{0}^{T}e^{-\lambda t}\left(t^{-\frac{1}{2}}+t^{-1+\alpha}\right)dt\to 0\quad\text{as }\lambda\to\infty.$
Thus, choosing $\lambda>0$ sufficiently large, we may assume that
$\|\kappa_{\lambda}\|_{L^{1}(0,T)}<1.$
By Young’s inequality for time convolution, we get
$\|F\|_{L^{2}(0,T)}\leq\|G\|_{L^{2}(0,T)}+\|\kappa_{\lambda}\star_{t}F\|_{L^{2}(0,T)}\leq\|G\|_{L^{2}(0,T)}+\|\kappa_{\lambda}\|_{L^{1}(0,T)}\|F\|_{L^{2}(0,T)}.$
It follows that
$\bigl(1-\|\kappa_{\lambda}\|_{L^{1}(0,T)}\bigr)\|F\|_{L^{2}(0,T)}\leq\|G\|_{L^{2}(0,T)},$
@@
$\|F\|_{L^{2}(0,T)}\leq\frac{1}{1-\|\kappa_{\lambda}\|_{L^{1}(0,T)}}\|G\|_{L^{2}(0,T)}.$
Finally, since $\|G\|_{L^{2}(0,T)}\leq\|g\|_{L^{2}(0,T)}$ and $\|f\|_{L^{2}(0,T)}\leq e^{\lambda T}\|F\|_{L^{2}(0,T)}$, we have
$\|f\|_{L^{2}(0,T)}\leq\frac{e^{\lambda T}}{1-\|\kappa_{\lambda}\|_{L^{1}(0,T)}}\|g\|_{L^{2}(0,T)}.$
This concludes the desired result. ∎
We now define
$g(t)$ $:=C\|\rho_{0}^{\varepsilon}-\rho_{0}\|_{H^{-\beta}}+C\left(\mathscr{H}[f^{\varepsilon}|M_{\rho,0}](0)+\mathscr{P}[\rho^{\varepsilon}|\rho](0)+\varepsilon^{2}\right)$
$\quad+C\left(\int_{0}^{t}(t-s)^{-\frac{1}{2}}\left(\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\frac{1}{f^{\varepsilon}}|(\xi-u^{\varepsilon})f^{\varepsilon}+\nabla_{\xi}f^{\varepsilon}|^{2}(s)\,dxd\xi\right)ds\right)^{\frac{1}{2}}.$
Then the inequality (4.9) takes the form
$\|(\rho^{\varepsilon}-\rho)(t)\|_{H^{-\beta}}\leq g(t)+C\int_{0}^{t}\bigl((t-s)^{-\frac{1}{2}}+(t-s)^{-1+\alpha}\bigr)\|(\rho^{\varepsilon}-\rho)(s)\|_{H^{-\beta}}\,ds.$
Moreover, $g\in L^{2}(0,T)$, and by Fubini’s theorem,
$\int_{0}^{T}|g(t)|^{2}\,dt$ $\leq C\left(\|\rho_{0}^{\varepsilon}-\rho_{0}\|_{H^{-\beta}}+\mathscr{H}[f^{\varepsilon}|M_{\rho,0}](0)+\mathscr{P}[\rho^{\varepsilon}|\rho](0)+\varepsilon^{2}\right)^{2}$
$\quad+C\int_{0}^{T}\int_{0}^{t}(t-s)^{-\frac{1}{2}}\left(\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\frac{1}{f^{\varepsilon}}|(\xi-u^{\varepsilon})f^{\varepsilon}+\nabla_{\xi}f^{\varepsilon}|^{2}(s)\,dxd\xi\right)dsdt$
$=C\left(\|\rho_{0}^{\varepsilon}-\rho_{0}\|_{H^{-\beta}}+\mathscr{H}[f^{\varepsilon}|M_{\rho,0}](0)+\mathscr{P}[\rho^{\varepsilon}|\rho](0)+\varepsilon^{2}\right)^{2}$
$\quad+C\int_{0}^{T}\left(\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\frac{1}{f^{\varepsilon}}|(\xi-u^{\varepsilon})f^{\varepsilon}+\nabla_{\xi}f^{\varepsilon}|^{2}(s)\,dxd\xi\right)\left(\int_{s}^{T}(t-s)^{-\frac{1}{2}}\,dt\right)ds$
$\leq C\left(\|\rho_{0}^{\varepsilon}-\rho_{0}\|_{H^{-\beta}}+\mathscr{H}[f^{\varepsilon}|M_{\rho,0}](0)+\mathscr{P}[\rho^{\varepsilon}|\rho](0)+\varepsilon^{2}\right)^{2}$
$\quad+C\int_{0}^{T}\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\frac{1}{f^{\varepsilon}}|(\xi-u^{\varepsilon})f^{\varepsilon}+\nabla_{\xi}f^{\varepsilon}|^{2}\,dxd\xi dt.$
Applying Lemma 4.2, we deduce
$\int_{0}^{T}\|(\rho^{\varepsilon}-\rho)(t)\|_{H^{-\beta}}^{2}\,dt$ $\leq C\left(\|\rho_{0}^{\varepsilon}-\rho_{0}\|_{H^{-\beta}}+\mathscr{H}[f^{\varepsilon}|M_{\rho,0}](0)+\mathscr{P}[\rho^{\varepsilon}|\rho](0)+\varepsilon^{2}\right)^{2}$
$\quad+C\int_{0}^{T}\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\frac{1}{f^{\varepsilon}}|(\xi-u^{\varepsilon})f^{\varepsilon}+\nabla_{\xi}f^{\varepsilon}|^{2}\,dxd\xi dt$
$\leq C\left(\|\rho_{0}^{\varepsilon}-\rho_{0}\|_{H^{-\beta}}+\mathscr{H}[f^{\varepsilon}|M_{\rho,0}](0)+\mathscr{P}[\rho^{\varepsilon}|\rho](0)+\varepsilon^{2}\right)^{2}$
due to (4.7). This completes the proof.
Remark 4.1.
A corresponding rigorous estimate toward the shifted Maxwellian $M_{\rho,{\rm u}_{\varepsilon}}$ can also be obtained in a negative Sobolev topology. Indeed, if we choose $\sigma>\max\{d,\beta\}$, then Corollary 3.1, combined with Theorem 1.1, implies
$\int_{0}^{T}\|f^{\varepsilon}-M_{\rho,{\rm u}_{\varepsilon}}\|_{H_{x,\xi}^{-\sigma}}^{2}\,dt$ $\leq C\int_{0}^{T}\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\frac{1}{f^{\varepsilon}}|\nabla_{\xi}f^{\varepsilon}+(\xi-u^{\varepsilon})f^{\varepsilon}|^{2}\,dxd\xi dt+C\int_{0}^{T}\int_{\mathbb{R}^{d}}\rho^{\varepsilon}|u^{\varepsilon}-{\rm u}_{\varepsilon}|^{2}\,dxdt$
$\quad+C\|{\rm u}_{\varepsilon}\|_{L^{\infty}((0,T)\times\mathbb{R}^{d})}^{2}\int_{0}^{T}\|\rho^{\varepsilon}-\rho\|_{L^{1}}^{2}\,dt+C\int_{0}^{T}\|\rho^{\varepsilon}-\rho\|_{H^{-\sigma}}^{2}\,dt$
$\leq C\left(\|\rho_{0}^{\varepsilon}-\rho_{0}\|_{H^{-\beta}}+\mathscr{H}[f^{\varepsilon}|M_{\rho,0}](0)+\mathscr{P}[\rho^{\varepsilon}|\rho](0)+\varepsilon^{2}\right)^{2},$
where we used $\|\rho^{\varepsilon}-\rho\|_{H^{-\sigma}}\leq\|\rho^{\varepsilon}-\rho\|_{H^{-\beta}}$.